The driver of a car suddenly applies the brakes hard for 2.6 s. As a result, an average force of 6200 N is exerted on the car to slow it. What is the impulse on the car?
1
Expert's answer
2020-03-19T11:29:46-0400

According to the second Newton's low


"F=\frac{\Delta p}{\Delta t}\to\Delta p=F\cdot\Delta t=6200\cdot2.6=16120" "\frac{kg\cdot m}{s}"